OpenStudy (anonymous):

what is the sqaure root of (x + 2y) = 10?

OpenStudy (mysesshou):

\[\sqrt{x+2y}=10\] I'm not sure which you're solving for. ?

OpenStudy (anonymous):

what is the value of x?

OpenStudy (mysesshou):

\[\sqrt{x + 2y}=10\] Ok. square both sides \[(\sqrt{x + 2y})^2=10^2\] \[x + 2y=100\] subtract 2y both sides \[x =100 -2y\]